OpenStudy (anonymous):

how calculate height of trapezoid without area? Or how calculate the area if I have B b and two diagonals?

OpenStudy (anonymous):

base and two diagonals?

OpenStudy (anonymous):

I have two bases and two diagonals? what formula I can use to calculate the area?

OpenStudy (anonymous):

Can you find the height based on the given information?

OpenStudy (anonymous):

Find the area of trapezoid with bases 7cm and 20cm and diagonals 13cm and 5\[\sqrt{10}\]

OpenStudy (anonymous):

this is all information which I have in problem

OpenStudy (anonymous):

One more hint, do you know what is the median of a trapezoid?

OpenStudy (anonymous):

yes, of gourse (B+b)/2; (20+7)/2=13.5cm

OpenStudy (anonymous):

I think I have a solution but it's a bit ugly, #1 Use this ( http://en.wikipedia.org/wiki/Trapezoid#Properties ) formula to find the the other two non-parallel sides of the trapezoid. #2 Find the height using this \[ h= \frac{\sqrt{(-a+b+c+d)(a-b+c+d)(a-b+c-d)(a-b-c+d)}}{2(b-a)} \] \( a\) and \( b\) are parallel sides and \(c\) and \( d\) are non-parallel sides. #3 Now you can find the area.

OpenStudy (anonymous):

But if it's given that that the diagonals intersects at right angle then you can simply find the area as \(\frac 12 \times\) product of the diagonals

OpenStudy (anonymous):

Thanks you much FoolForMath. But without length of non-parallel sides, or without height I can't find area. Maybe is another way to solve.